# Variational Learning of Disentangled Representations

### v1

@@ A.2.1 Enviroment balancing methods
The field of out-of-distribution (OOD) generalization emerged from foundational work on causality and invariance across training environments [Peters et al., 2016, 2017]. The central assumption is that each environment exhibits distinct spurious correlations between features and labels; therefore, robust generalization requires models to focus on invariant relationships that hold across environments. To address this distribution shift, many recent approaches adopt a regularized empirical risk minimization framework:
@@
where the regularizer $R$ encourages representations that are stable across environments. Among these, Invariant Risk Minimization (IRM) [Arjovsky et al., 2019] enforces that a single classifier remains optimal across all environments, Variance Risk Extrapolation (VarREx) [Krueger et al., 2021] promotes robustness by minimizing the variance of losses across environments, and CLOvE [Wald et al., 2021] takes a calibration-theoretic perspective, penalizing discrepancies between predicted confidence and correctness across environments.
@@ A.2.2 Distributionally robust methods
An alternative line of work for handling distribution shifts is Distributionally Robust Optimization (DRO) [Ben-Tal et al., 2013, Duchi et al., 2021, Duchi and Namkoong, 2021, Wei et al., 2023], which avoids assuming a fixed data-generating distribution. Instead, DRO methods optimize performance under the worst-case scenario over a family of plausible distributions. A prominent variant, known as group DRO [Sagawa et al., 2019, Piratla et al., 2021], introduces group-level structure that may correlate with spurious features, potentially leading to biased predictions. In settings where group labels are not directly observed, several strategies have been proposed, including reweighting high-loss examples [Liu et al., 2021] and balancing class-group combinations through data sub-sampling [Idrissi et al., 2022].
@@ A.3 Zero-shot learning
Zero-shot learning systems [Fei-Fei et al., 2006, Larochelle et al., 2008] aim to classify instances from novel, previously unseen classes at test time. In contrast to the out-of-distribution (OOD) generalization setting, these approaches typically do not assume the presence or structure of a distribution shift. Instead, a common strategy is to learn data representations that capture class-agnostic similarity, enabling the model to determine whether two instances belong to the same class without requiring knowledge of the class identity itself. Such methods include contrastive-learning [Hadsell et al., 2006], siamese neural networks [Koch et al., 2015], triplet networks [Hoffer and Ailon, 2015], and other more recent variations [Oh Song et al., 2016, Sohn, 2016, Wu et al., 2017, Yuan et al., 2019]. Recent work has begun to address the impact of class distribution shifts in zero-shot settings. For instance, Slavutsky and Benjamini [2024] integrate environment-based regularization—motivated by OOD generalization—with zero-shot learning by simulating distribution shifts through hierarchical sampling, enabling the model to learn representations that are robust to shifts in class distributions.
@@ Appendix B Proofs
$=\left[\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}(q_{w|x,y}\,\|\,p_{w|x,y})\right]-\left[\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)\right]$ (14)
$=\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$ (15)
@@
$=\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)-\log p(z\mid w,x,y)\right]$ (19)
$=\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right).$ (20)
@@ B.2.1 Regularity conditions
To ensure that expectations and KL-terms in the game objective $\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})$ render the functionals strictly concave in $q_{z|x}$ , strictly concave in $q_{w|x,y}$, and strictly convex in $g$, the following regularity conditions are required:
@@
    The variational families $Q_{z}$ and $Q_{w}$, and the set of achievable classifiers $\mathcal{G}$ are non-empty, convex and compact.
@@
$\mathcal{L}_{z}(q_{z|x},p;x)=\log p(x)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right).$ (21)
@@
$\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ $=\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (22)
@@
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})=\mathbb{E}_{p_{x,y}}$ $\left[\log p(x)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right)\right.$ (23)
$+\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (24)
@@
$\ell_{w}\coloneqq$ $-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (27)
[…]
@@
$\ell_{z}$ $\coloneqq-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (32)
By convexity of KL divergence in its first argument, $-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right)$ is strictly concave in $q_{z\mid x}$.
@@
$-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)=H(q_{z\mid x})+H(q_{w\mid x,y})+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y),$ (33)
@@
where $-\left\|w-\mu_{y}\right\|^{2}$ is a quadratic form in $\mu_{y}$, which is linear in $q_{z|x}$, and thus $\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)\right]$ is strictly concave in $q_{z|x}$. Hence, $-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ is strictly concave in $q_{z|x}$, and thus so is $\mathbb{E}_{p_{x,y}}[\ell_{z}]$. ∎
@@ Appendix C Supplementary Figures
Appendix D Supplementary Tables
@@
Marginal RMSE ($\rho=0$) $\downarrow$
@@
DISCoVeR (Ours) $\mathbf{0.064\pm 0.000}$
@@
CCVAE $0.092\pm 0.002$ $0.103\pm 0.014$ $0.099\pm 0.011$ $0.092\pm 0.005$
DISCoVeR (Ours) $\mathbf{0.073\pm 0.001}$ $\mathbf{0.083\pm 0.004}$ $\mathbf{0.087\pm 0.002}$ $\mathbf{0.087\pm 0.001}$
@@
DISCoVeR (Ours) $\mathbf{0.030\pm 0.011}$ $135.677\pm 0.007$
DISCoVeR - Common (Ours) $374.114\pm 0.05$
Appendix E Impelementation Details
E.1 Considerations and reproducibility
We run all experiments on a single H100 GPU. Reported means and standard deviations for tables are conducted over 10 repetitions of the experiment with different random seeds. All models are trained using the AdamW [Loshchilov and Hutter, 2019] optimizer until validation loss stops decreasing for 50 epochs. Wherever provided, we use mutual information neural estimation (MINE, Belghazi et al. [2018]) to obtain mutual information estimates. For Naive Bayes classifiers, we use the implementation provided by scikit-learn [Pedregosa et al., 2011]. To use ideal hyperparameters for each method, we consult the original implementation whenever possible, and conduct a simple grid-search to produce originally described model behavior. Implementations of all methods compared in this study, including DISCoVeR, as well as code to reproduce our results, can be found at https://github.com/Computational-Morphogenomics-Group/DISCoVeR. Models compared in the study admit a weighting term for each term in the loss function, of which most are shared across different approaches. We use the following shorthands for each of the terms:
@@ E.1 Considerations and reproducibility
$\operatorname{D_{\mathrm{KL}}}(Z)\rightarrow\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z})$
$\operatorname{D_{\mathrm{KL}}}(W)\rightarrow\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right)$

### v2

@@ A.2.1 Enviroment balancing methods
The field of out-of-distribution (OOD) generalization emerged from foundational work on causality and invariance across training environments (peters2016causal; peters2017elements). The central assumption is that each environment exhibits distinct spurious correlations between features and labels; therefore, robust generalization requires models to focus on invariant relationships that hold across environments. To address this distribution shift, many recent approaches adopt a regularized empirical risk minimization framework:
@@
where the regularizer $R$ encourages representations that are stable across environments. Among these, Invariant Risk Minimization (IRM) (arjovsky2019invariant) enforces that a single classifier remains optimal across all environments, Variance Risk Extrapolation (VarREx) (krueger2021out) promotes robustness by minimizing the variance of losses across environments, and CLOvE (wald2021calibration) takes a calibration-theoretic perspective, penalizing discrepancies between predicted confidence and correctness across environments.
@@ A.2.2 Distributionally robust methods
An alternative line of work for handling distribution shifts is Distributionally Robust Optimization (DRO) (DRO; duchi2021statistics; duchi2021learning; wei2023distributionally), which avoids assuming a fixed data-generating distribution. Instead, DRO methods optimize performance under the worst-case scenario over a family of plausible distributions. A prominent variant, known as group DRO (sagawa2019distributionally; piratla2021focus), introduces group-level structure that may correlate with spurious features, potentially leading to biased predictions. In settings where group labels are not directly observed, several strategies have been proposed, including reweighting high-loss examples (liu2021just) and balancing class-group combinations through data sub-sampling (idrissi2022simple).
@@ A.3 Zero-shot learning
Zero-shot learning systems (few_shot; zero_shot) aim to classify instances from novel, previously unseen classes at test time. In contrast to the out-of-distribution (OOD) generalization setting, these approaches typically do not assume the presence or structure of a distribution shift. Instead, a common strategy is to learn data representations that capture class-agnostic similarity, enabling the model to determine whether two instances belong to the same class without requiring knowledge of the class identity itself. Such methods include contrastive-learning (contrastive), siamese neural networks (siamese_networks), triplet networks (triplet_networks), and other more recent variations (oh2016deep; sohn2016improved; wu2017sampling; yuan2019signal). Recent work has begun to address the impact of class distribution shifts in zero-shot settings. For instance, slavutsky2024class integrate environment-based regularization—motivated by OOD generalization—with zero-shot learning by simulating distribution shifts through hierarchical sampling, enabling the model to learn representations that are robust to shifts in class distributions.
@@ Appendix B Proofs
$=\left[\log p(x\mid y)-D_{\mathrm{KL}}(q_{w|x,y}\,\|\,p_{w|x,y})\right]-\left[\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)\right]$ (13)
$=D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)-D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$ (14)
@@
$=\mathbb{E}_{q_{w|x,y}}\left[\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)-\log p(z\mid w\mid x,y)\right]\right]$ (18)
$=\mathbb{E}_{q_{w|x,y}}\left[\text{KL}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right)\right].$ (19)
@@ B.2.1 Regularity conditions
To ensure that expectations and KL‑terms in the game objective $\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})$ render the functionals strictly concave in $q_{z|x}$ , strictly concave in $q_{w|x,y}$, and strictly convex in $g$, the following regularity conditions are required:
@@
    The variational families $Q_{z}$ and $Q_{w}$, and the set of achievable classifiers $\mathcal{G}$ are non‑empty, convex and compact.
@@
$\mathcal{L}_{z}(q_{z|x},p;x)=\log p(x)-D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right).$ (20)
@@
$\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ $=\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (21)
@@
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})=\mathbb{E}_{p_{x,y}}$ $\left[\log p(x)-D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right)\right$. (22)
$+\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (23)
@@
$\ell_{w}\coloneqq$ $-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (26)
$=$ $-\iint q\left(z\mid x\right)q(w\mid x,y)\left[\log q\left(z\mid x\right)+\log q(w\mid x,y)-\log p(z,w\mid x,y)\right]\,dz\,dw$
$=$ $-\int q\left(z\mid x\right)\log q\left(z\mid x\right)\,dz+\int q(w\mid x,y)\log q(w\mid x,y)\,dw$ (27)
$-\iint q\left(z\mid x\right)q(w\mid x,y)\log p(z,w\mid x,y)\,dz\,dw$ (28)
@@
$\ell_{z}$ $\coloneqq-D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (31)
By convexity of KL divergence in its first argument, $D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right)$ is strictly concave in $q_{z\mid x}$.
@@
$D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)=H(q_{z\mid x})+H(q_{w\mid x,y})+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y),$ (32)
@@
where $\left\|w-\mu_{y}\right\|^{2}$ is a quadratic form in $\mu_{y}$, which is linear in $q_{z|x}$, and thus $\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)\right]$ is strictly concave in $q_{z|x}$. Hence, $D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ is strictly concave in $q_{z|x}$, and thus so is $\mathbb{E}_{p_{x,y}}[\ell_{z}]$. ∎
@@ Appendix C Supplementary Figures
Appendix D Supplementary tables for experimental results
@@
Marginal RMSE ($p=0$) $\downarrow$
@@
DisCoVR (Ours) $\mathbf{0.064\pm 0.000}$
@@
CCVAE $0.092\pm 0.002$ $103\pm 0.014 0.0$$99\pm 0.011 0.0$$092\pm 0.005$
DisCoVR (Ours) $\mathbf{0.073\pm 0.001}$ $\mathbf{0.083\pm 0.004}$ $\mathbf{0.087\pm 0.002}$ $\mathbf{0.087\pm 0.001}$
@@
DisCoVR (Ours) $\mathbf{0.030\pm 0.011}$ $135.677\pm 0.007$
DisCoVR - Common (Ours) $374.114\pm 0.05$
Appendix E Additional disentanglement metrics
We provide an extended disentanglement assessment using multiple metrics. Because mutual information is difficult to estimate reliably, we report two estimators—MINE and kNN. Although their absolute values differ, the relative rankings of the methods remain consistent as can be seen in the ranking tables. In addition to these mutual-information estimates, we also report the following metrics, which quantifying the level of label information captured by $w$ compared to $z$ :
Mutual Information Gap (MIG)
$\operatorname{MIG}(w;z)=\frac{I(y;w)-I(y;z)}{H(y)}$
Mutual Information Completeness (MIC)
$\operatorname{MIC}(w;z)=\frac{I(y;w)}{I(y;w)+I(y;z)}$
E.1 Parametric model
CSVAE and its variants impose a fully separable prior, thereby forcing separability even when the true latent structure is not separable (see Table 1). In contrast, DisCoVR learns informative conditional embeddings that closely track the true posterior without requiring ground-truth knowledge of a truncated or fully separable prior, and it outperforms both DIVA and CCVAE.
Replacing the prior in DisCoVR with a fully separable predefined prior on $w$ yields consistent embeddings with the ground-truth structure while retaining the benefits of separability.
Supplementary Table 4: Additional disentanglement metrics calculated with kNN mutual information estimation for the parametric model dataset with $k=20$. Bold indicates closest to true posterior within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
@@ E.1 Parametric model
[…]
HCSVAE - N.A. $0.000\pm 0.000$ $\mathbf{0.643\pm 0.000}$ $000\pm 0.001 0.0$$072\pm 0.000$ $1.000\pm 0.000$ $0.000\pm 0.000$
HCSVAE $0.000\pm 0.000$ $\mathbf{0.643\pm 0.001}$ $001\pm 0.001 0.0$$072\pm 0.000$ $1.000\pm 0.000$ $0.000\pm 0.001$
DisCoVR (CSVAE prior) $0.000\pm 0.000$ $\mathbf{0.643\pm 0.000}$ $051\pm 0.007 0.0$$072\pm 0.000$ $1.000\pm 0.000$ $\mathbf{0.031\pm 0.005}$
Flexible DIVA $021\pm 0.042 0.0$$91\pm 0.046 0.0$$00\pm 0.000 0.0$$008\pm 0.010$ $800\pm 0.400 0.0$$000\pm 0.000$
CCVAE $\mathbf{0.022\pm 0.043}$ $90\pm 0.045 0.000$$0\pm 0.000 0.0$$008\pm 0.010$ $800\pm 0.400 0.0$$000\pm 0.000$
DisCoVR (our prior) $0.010\pm 0.006$ $\mathbf{0.151\pm 0.007}$ $\mathbf{0.108\pm 0.029}$ $\mathbf{0.016\pm 0.001}$ $\mathbf{0.938\pm 0.035}$ $\mathbf{0.072\pm 0.020}$
Fully Separable Posterior (no truncation) $057\pm 0.001 0.0$$057\pm 0.000$ $144\pm 0.003 0.0$$000\pm 0.000$ $499\pm 0.006 0.0$$090\pm 0.002$
True Posterior $0.058\pm 0.003$ $0.643\pm 0.000$ $144\pm 0.005 0.0$$066\pm 0.000$ $917\pm 0.004 0.0$$055\pm 0.003$
Supplementary Table 5: Additional disentanglement metrics calculated with MINE mutual information estimation for the parametric model dataset. Bold indicates closest to true posterior within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
CSVAE $0.033\pm 0.055$ $\mathbf{0.526\pm 0.045}$ $031\pm 0.045 0.0$$055\pm 0.010$ $945\pm 0.091 0.0$$009\pm 0.005$
HCSVAE - N.A. $0.000\pm 0.000$ $543\pm 0.018 0.0$$00\pm 0.000 0.0$$061\pm 0.002$ $1.000\pm 0.000$ $0.000\pm 0.000$
HCSVAE $0.000\pm 0.000$ $543\pm 0.022 0.0$$00\pm 0.000 0.0$$061\pm 0.002$ $1.000\pm 0.000$ $0.001\pm 0.000$
[…]
Flexible DIVA $0.027\pm 0.053$ $113\pm 0.056 0.0$$01\pm 0.001 0.0$$010\pm 0.012$ $798\pm 0.398 0.0$$001\pm 0.000$
CCVAE $0.026\pm 0.052$ $115\pm 0.058 0.0$$01\pm 0.001 0.0$$010\pm 0.012$ $799\pm 0.399 0.0$$001\pm 0.000$
DisCoVR (ours) $\mathbf{0.037\pm 0.006}$ $\mathbf{0.176\pm 0.008}$ $\mathbf{0.109\pm 0.025}$ $\mathbf{0.016\pm 0.001}$ $\mathbf{0.825\pm 0.026}$ $\mathbf{0.073\pm 0.018}$
Fully Separable Posterior (no truncation) $084\pm 0.003 0.0$$083\pm 0.003$ $137\pm 0.004 0.0$$000\pm 0.000$ $497\pm 0.009 0.0$$088\pm 0.003$
True Posterior $0.085\pm 0.005$ $0.493\pm 0.011$ $139\pm 0.005 0.0$$046\pm 0.002$ $853\pm 0.009 0.0$$057\pm 0.004$
Supplementary Table 6: Rank (1 = closest to True Posterior) of each method with respect to the true posterior for metrics calculated with kNN mutual information estimation with $k=20$. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
@@
[…]
Supplementary Table 7: Rank (1 = closest to True Posterior) of each method with respect to the True Posterior for metrics calculated with MINE mutual information estimation. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Fully Separable DisCoVR (CSVAE prior) $cellcolorUpper!55!Lower 3 \cellcolorUpper$!70!Lower 2 $cellcolorUpper!55!Lower 3 \cellcolorUpper$!85!Lower 1
[…]
E.2 Noisy Swiss Roll
When the observed labels are noisy, DisCoVR outperforms other methods, obtaining embeddings close to the ground truth.
Supplementary Table 8: Additional disentanglement metrics calculated with kNN mutual information estimation for the Noisy Swiss Roll ($p=0.3$) dataset with $k=20$. Bold indicates closest to ground truth within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
Fully Separable CSVAE - N.A. $0.041\pm 0.007$ $0.525\pm 0.221$ $362\pm 0.180 0.0$$057\pm 0.026$ $0.888\pm 0.098$ $0.266\pm 0.152$
[…]
HCSVAE - N.A. $0.029\pm 0.007$ $642\pm 0.000 0.0$$65\pm 0.013 0.0$$072\pm 0.001$ $957\pm 0.010 0.0$$009\pm 0.019$
[…]
Flexible DIVA $034\pm 0.013 0.0$$036\pm 0.011$ $2.633\pm 0.360$ $0.000\pm 0.003$ $0.515\pm 0.159$ $2.185\pm 0.332$
CCVAE $040\pm 0.015 0.0$$030\pm 0.007$ $2.952\pm 0.124$ $0.001\pm 0.003$ $0.447\pm 0.153$ $2.462\pm 0.118$
DisCoVR (ours) $\mathbf{0.000\pm 0.000}$ $\mathbf{0.049\pm 0.002}$ $\mathbf{0.029\pm 0.011}$ $\mathbf{0.006\pm 0.000}$ $\mathbf{1.000\pm 0.000}$ $\mathbf{0.014\pm 0.008}$
Noisy Ground Truth $000\pm 0.000 0.0$$55\pm 0.002 0.0$$00\pm 0.000 0.0$$007\pm 0.000$ $1.000\pm 0.000$ $0.000\pm 0.000$
Supplementary Table 9: Additional disentanglement metrics calculated with MINE mutual information estimation for the Noisy Swiss Roll ($p=0.3$) dataset. Bold indicates closest to ground truth within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
CSVAE - N.A. $0.046\pm 0.022$ $422\pm 0.186 0.0$$50\pm 0.020 0.0$$044\pm 0.023$ $834\pm 0.184 0.0$$029\pm 0.017$
[…]
HCSVAE - N.A. $0.023\pm 0.014$ $585\pm 0.011 0.0$$26\pm 0.012 0.0$$066\pm 0.002$ $963\pm 0.021 0.0$$006\pm 0.002$
[…]
Noisy Ground Truth $000\pm 0.000 0.0$$24\pm 0.018 0.0$$00\pm 0.001 0.0$$003\pm 0.002$ $985\pm 0.048 0.0$$002\pm 0.001$
Supplementary Table 10: Rank (1 = closest to Ground Truth) of each method with respect to the Ground Truth for metrics calculated with kNN mutual information estimation with $k=20$ on the Noisy Swiss Roll ($p=0.3$) dataset. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Method $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Supplementary Table 11: Rank (1 = closest to Ground Truth) of each method with respect to the Ground Truth for metrics calculated with MINE mutual information estimation on the Noisy Swiss Roll ($p=0.3$) dataset. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Appendix F Ablations on model components
We evaluate the contribution of each model component by examining two variations: (1) applying the classifier directly to $z$, and (2) replacing the conditional prior on $w$ with a standard Gaussian.
When training the classifier directly on $z$ we were able achieve results qualitatively similar to those obtained using the reconstruction $\hat{x}$, but doing so requires substantially more parameter tuning.
An unconditional standard Gaussian prior for $w$, causes $w$ to collapse into a representation redundant with $z$, removing meaningful separation.
Supplementary Figure 7: Ablation study on the Noisy Swiss Roll ($p=0.3$) dataset.
Supplementary Figure 8: Ablation study on the Noisy Colored MNIST ($p=0.3$) dataset. For each setting: left column denotes coloring by noisy labels, right column denotes coloring by digit (shape, not included in the label).
Appendix G Additional experiment on CelebA-Hats
We performed an additional experiment on the CelebA dataset, with the attribute Wearing_hatWearing\_hat denoting the $y$ label. Supplementary Table 12 outlines the results of this experiment. DisCoVR is the only method that exhibits high disentanglement for $z,w$ without compromising reconstruction quality.
Supplementary Table 12: Model performances of a single experiment on CelebA-Hats. Bold denotes best performance.
$I(z;w)\downarrow$ NLL ($\downarrow$)
CSVAE - N.A. 0.360 653.537
CSVAE 0.213 351.082
HCSVAE - N.A. 0.135 2608.442
HCSVAE 0.192 673.674
DIVA 0.553 356.090
CCVAE 0.856 347.940
DisCoVR (Ours) 0.059 353.271
DisCoVR (Ours) - Common - 437.144
Appendix H Implementation details
H.1 Considerations and reproducibility
We run all experiments on a single H100 GPU. Reported means and standard deviations for tables are conducted over 10 repetitions of the experiment with different random seeds. All models are trained using the AdamW (loshchilov2018decoupled) optimizer until validation loss stops decreasing for 50 epochs. Wherever provided, we use mutual information neural estimation (MINE, pmlr-v80-belghazi18a) and k-Nearest Neighbor (kNN) mutual information estimation Kraskov2011 to obtain mutual information estimates. For Naive Bayes classifiers, we use the implementation provided by scikit-learn (scikit-learn). To use ideal hyperparameters for each method, we consult the original implementation whenever possible, and conduct a simple grid-search to produce originally described model behavior. Implementations of all methods compared in this study, including DisCoVR, as well as code to reproduce our results, is attached to this submission and will be made public upon acceptance. Models compared in the study admit a weighting term for each term in the loss function, of which most are shared across different approaches. We use the following shorthands for each of the terms:
$\text{Rec.}\rightarrow\mathbb{E}_{q_{z|x}}\left[\mathbb{E}_{q_{w|x,y}}[\log p(x\mid z,w)]\right]$
$D_{\mathrm{KL}}(Z)\rightarrow D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z})$
$D_{\mathrm{KL}}(W)\rightarrow D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right)$
$\text{Adv.}\rightarrow-\mathbb{E}_{q_{z|x}}[\log g(y\mid z)]$
$\text{Class.}\rightarrow\mathbb{E}_{q_{w|x,y}}[\log q(y\mid w)]$
$\text{Rec. - ($Z$)}\rightarrow\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]$